OpenStudy (anonymous):

what is 1/2 plus 0.25 equil?

OpenStudy (anonymous):

1/2=.5 so 1/2+.25=.75

OpenStudy (anonymous):

.75 and 3/4 are both correct

OpenStudy (mathstudent55):

\(\dfrac{1}{2} + 0.25 \) \( = 0.5 + 0.25 \) Now you can finish it easily.

OpenStudy (mathstudent55):

If you prefer to do it with fractions, then: \(\dfrac{1}{2} + 0.25 \) \( = \dfrac{1}{2} + \dfrac{1}{4} \) \(= \dfrac{1}{2} \times \dfrac{2}{2} + \dfrac{1}{4} \) \( = \dfrac{2}{4} + \dfrac{1}{4} \) \(= \dfrac{3}{4} \)